A metal occurs as body centered cube and has a density of 7.856 g/cm3. calculate the atomic radius of metal. atomic mass of metal is 58 g/ mol. find atomic radius of metal. 

1
Expert's answer
2021-11-12T08:16:01-0500


a= "\frac{4r}{√m}"

7.856= 4×r/√58

7.856= 4r/7.62

4r= 7.62×7.856

4r= 59.86

r=14.97cm